# Equilibrium of Charges and Differential Equations Solved by Polynomials II

### v2

@@ 1 Introduction
System (1) is invariant under rigid motions and scaling of the plane. From invariance of the stationary value of the coulomb energy under scaling of the plane it follows that the system has solutions only if $\sum_{i<j}Q_{i}Q_{j}=0$. O’Neil showed [25] that for almost all values of $Q_{i}$, satisfying the above condition, there are exactly $(N-2)!$ distinct configurations, i.e. distinct, modulo rigid motions and scaling of the plane, solutions of (1). However, there are “resonant” values of $Q_{i}$, where number of configurations is infinite.
@@
where prime denotes differentiation wrt $z$ and
@@ 2 Darboux Transformations of Second Order Operators and Equilibrium of Charges
where $A$ and $A^{*}$ are (formally conjugated) first order differential operators
@@
where $C$ is an arbitrary constant, and the primitive of $\varkappa^{2}$ includes another arbitrary constant of integration. This extends the Darboux transformation (14) of an eigenfunction to the case $\lambda=\eta$. Transformation (15) is usually considered for $\eta=0$ and is called the Darboux transformation at zero energy level.
@@
$\left(-\partial^{2}-2(\log q)^{\prime\prime}\right)\left[\frac{p}{q}\right]=0,$ (16)
@@
we get the first-order recurrence relation for the Adler-Moser polynomials (6). Therefore, given the initial conditions $P_{0}=1$, $P_{1}=z$, the Darboux transformation (15,17) generates all possible equilibrium configurations of charges with $Q_{i}\in\{-1,1\}$.
@@ 3 $\Lambda=2$ Case and Third-Order Operators
However, we cannot apply Darboux transformations to the above second order operators, because now $u_{n}=-6(\log q_{n})^{\prime\prime}$ (and not $-2(\log q_{n})^{\prime\prime}$, cf. (13,16)). To proceed, we note that according to the Abel identities (8), $p_{n}$ in (21, 22) can be replaced with $q_{n+1}^{\prime}q_{n}-q_{n+1}q_{n}^{\prime}$ and we get the third-order differential equation
@@ 4 Darboux Transformations for Third-Order Operators
were found by Aiyer et al [4]. Their derivation, via factorization, was presented by Athorne and Nimmo in [5]. Similarly to the $\Lambda=1$ case, we need their “zero-level” extensions (for $\lambda=\eta=0$) which we will find in this section.
@@
In contrast to the case of the Schrodinger operator, permutation of factors in (25) maps $L$ into $\hat{L}$ of the similar type (i.e. as in (24)) just in the trivial case $u={\rm const}$. Non-trivial result arises from the transformation
@@
Equation (28) has a solution $\hat{f}=-f$ and the Darboux-type transform $\hat{L}=\hat{B}\hat{A}$ is obtained from $L=BA$ by the involution [^5] This is similar to the Schrodinger operator case where the Darboux transformation is the involution $A=\partial-f\leftrightarrow A^{*}=-(\partial+f)$, cf (11). $f\to-f$. From (25, 26) it follows that for such a transformation
@@
where the primitives include arbitrary constants of integration. Thus, we have found the zero-level Darboux transformations (29, 32) for the third-order operators of the type $\partial^{3}-u\partial$.
@@ 5 $\Lambda=2$ Configurations and Darboux Transformations
Substituting (33,34) into transformation (32), or equivalently, substituting $\hat{\phi}=p/q^{2}$, $(\varkappa^{-1})^{\prime}=-\varkappa^{-2}p_{-}/q_{-}^{2}$ and $\varkappa=q/q_{-}$ into (31) and rewriting result in the differential form, we obtain
@@
Equations (35, 36) are, modulo constant factors, the first-order recurrence relations (8) with $q_{+}=q_{n+1}$, $q=q_{n}$, $q_{-}=q_{n-1}$ and $p=p_{n}$, $p_{-}=p_{n-1}$. Together with initial conditions $q_{0}=p_{0}=1$, they produce solutions of (7). There also exist other initial conditions for which solutions of (8) are polynomials that are not solutions of (7). They will be considered in the next section.
@@ 6 Darboux Transformations and Terminating Configurations
As was mentioned in section 2, the first-order recurrence relation (6) for $P_{n}$ follows from the Darboux transformation (15, 17). Bilinear equation (5) is not involved in this derivation. Therefore, to get a sequence of equilibrium configurations, the initial conditions for (6) need not necessarily to be those for the Adler-Moser polynomials $P_{0}=1$, $P_{1}=z$. A family of sequences with other initial conditions was introduced in work by Duistermmat and Grunbaum [12] in the context of bi-spectral problem. In relation to vortex equilibria this family appeared in work by O’Neil and Cox-Steib [27] (also see [18]). There, sequences start from $P_{0}=1$, $P_{1}=z^{1/2}$. In difference from the Adler-Moser case, these sequences terminate as a non-factorizable $\varkappa$ emerges at some step of recurrence. Examples of first several sequences are:
@@
Similar situation occurs in the case of the third-order operators. As we saw in section 5, the Darboux transformations (32, 37), together with (38), result in the first-order recurrence relations (8). For two species of charges initial conditions are $p_{0}=q_{0}=1$. However, there exist other initial conditions, where terminating sequences of equilibrium configurations, that are not solutions of (7), are generated by the Darboux transformations. Such sequences of solutions were found in [27] (also see [10], [18]). For example, for $q_{0}=1$, $p_{0}=z^{2}$ the first several sequences are
@@ 7 Conclusions and Open Problems
When $\Lambda=1$, this generalization is associated with the soliton solutions of the KdV hierarchy and “trigonometric” Baker-Akhieser function $\Phi=q(z,k)/p(z,k)e^{kz}$, where now $q$ and $p$ are trigonometric polynomials in $z$ ([14, 17, 21], for review see e.g. [22]). Trigonometric Baker-Akhiezer function is obtained from the eigenfunction $e^{kz}$ (with eigenvalue $\lambda=k^{2}$) of the free Schrodinger operator $\partial^{2}$ by iterations of the Darboux transformations: First with $\eta=k_{1}^{2}$, then transformation with $\eta=k_{2}^{2}$, etc. Here $k_{i}$ are integers, such that $0<k_{1}<k_{2}<\dots$.
When $\Lambda=2$, translating vortex streets configurations cannot be generated by Darboux transformations: Even in the non-translating case $k=0$, only short terminating sequences can be produced by the zero-level transformations [22]. Indeed, similarly to the second order case, operators $L$ with non-zero periodic coefficients can be obtained by sequence of Darboux transformations from $\partial^{3}$ only when transformations are associated with a sequence of distinct, nonzero $\eta$. Here equations
$L_{i}\varkappa_{i}=\eta_{i}\varkappa_{i},\quad L_{i}=\partial^{3}-u_{i}\partial,\quad\eta_{i}=k_{i}^{3},$
where $u_{i}$ is periodic and $k_{i}\not=0$, cannot be reduced to the Schrodinger equations. As a consequence, $\varkappa_{i}$ do not correspond to equilibrium configurations. It would be interesting to explore the $\Lambda=2$ vortex street configurations using alternative methods.

### v4

@@ 1 Introduction
System (1) is invariant under rigid motions and scaling of the plane. From invariance of the stationary value of the coulomb energy under scaling of the plane it follows that the system has solutions only if $\sum_{i<j}Q_{i}Q_{j}=0$. O’Neil showed [25] that for almost all values of $Q_{i}$, satisfying the above condition, there are exactly $(N-2)!$ distinct configurations, i.e. distinct, modulo translations, rotations and scaling of the plane, solutions of (1). However, there are “resonant” values of $Q_{i}$, where number of configurations is infinite.
@@
where prime denotes differentiation wrt $z$ and polynomials
@@ 2 Darboux Transformations of Second Order Operators and Equilibrium of Charges
where $A$ and $A^{*}$ are (formally adjoint) first order differential operators
@@
where $C$ is an arbitrary constant, and the primitive of $\varkappa^{2}$ includes another arbitrary constant of integration. This extends the Darboux transformation (14) of an eigenfunction to the case $\lambda=\eta$. Iterations of transformation (15) are typically considered for $\eta=0$ and are called the Darboux transformations at zero energy level.
@@
$\left(-\partial^{2}-2(\log q)^{\prime\prime}\right)\left[\frac{p}{q}\right]=0,$
@@
we get the first-order recurrence relation for the Adler-Moser polynomials (6). Therefore, given the initial conditions $P_{0}=1$, $P_{1}=z$, the Darboux transformation (15,16) generates all possible equilibrium configurations of charges with $Q_{i}\in\{-1,1\}$.
@@ 3 $\Lambda=2$ Case and Third-Order Operators
However, sequence of the above second-order operators cannot be generated by Darboux transformations (13), because now $u_{n\pm 1}\not=u_{n}-2(\log\phi)^{\prime\prime}$. To proceed, we note that according to the Abel identities (8), $p_{n}$ in (20, 21) can be replaced with $q_{n+1}^{\prime}q_{n}-q_{n+1}q_{n}^{\prime}$ and we get the third-order differential equation
@@ 4 Darboux Transformations for Third-Order Operators
were found by Aiyer et al [4]. Their derivation, via factorization, was presented by Athorne and Nimmo in [5]. Similarly to the $\Lambda=1$ case, we need “zero-level” transformations (for $\lambda=\eta=0$) which we will find in this section.
@@
In contrast to the case of the Schrodinger operator, permutation of factors in (24) maps $L$ into $\hat{L}$ of the similar type – that is, an operator without the zeroth-order term, as in (23) – only if [^5] This corresponds to $u={a}^{2}\left(6\,c\,{\rm cn}\left(az+b,c\right){\rm dn}\left(az+b,c\right)+6\,{c}^{2}\,{\rm sn}^{2}\left(az+b,c\right)-{c}^{2}-1\right)$ and $\hat{u}=a^{2}(6\,{c}^{2}{\rm sn}^{2}\left(az+b,c\right)-{c}^{2}-1)$. In particular, in the rational limit, $u=0$ or $u=12/z^{2}$, where $v=2/z$ or $v=-2/z$ respectively. $v=2ac\,{\rm sn}(az+b,c)$, where $a,b$ and $c$ are arbitrary constants. In general, to obtain an operator of the same type, several permutations and re-factorizations are required (see [22] for more details). Instead of studying these intermediate operations separately, we consider the transformation
@@
Equation (27) has a solution $\hat{f}=-f$ and the Darboux-type transform $\hat{L}=\hat{B}\hat{A}$ is obtained from $L=BA$ by the involution [^6] This is similar to the Schrodinger operator case where the Darboux transformation is the involution $A=\partial-f\leftrightarrow A^{*}=-(\partial+f)$, cf (11). $f\to-f$. From (24, 25) it follows that for such a transformation
@@
where the primitives include arbitrary constants of integration. Thus, we have found the zero-level Darboux transformations (28, 31) for the third-order operators of the type $\partial^{3}-u\partial$.
@@ 5 $\Lambda=2$ Configurations and Darboux Transformations
Substituting (32,33) into transformation (31), or equivalently, substituting $\hat{\phi}=p/q^{2}$, $(\varkappa^{-1})^{\prime}=-\varkappa^{-2}p_{-}/q_{-}^{2}$ and $\varkappa=q/q_{-}$ into (30) and rewriting result in the differential form, we obtain
@@
Equations (34, 35) are, modulo nonessential constant factors, the first-order recurrence relations (8) with $q_{+}=q_{n+1}$, $q=q_{n}$, $q_{-}=q_{n-1}$ and $p=p_{n}$, $p_{-}=p_{n-1}$. Together with initial conditions $q_{0}=p_{0}=1$, they produce solutions of (7). There also exist other initial conditions for which solutions of (8) are polynomials that are not solutions of (7). They will be considered in the next section.
@@ 6 Darboux Transformations and Terminating Configurations
As was mentioned in section 2, the first-order recurrence relation (6) for $P_{n}$ follows from the Darboux transformation (15, 16). Bilinear equation (5) is not involved in this derivation. Therefore, to get a sequence of equilibrium configurations, the initial conditions for (6) need not necessarily to be those for the Adler-Moser polynomials $P_{0}=1$, $P_{1}=z$. A family of sequences with other initial conditions was introduced in work by Duistermmat and Grunbaum [12] in the context of bi-spectral problem. In relation to vortex equilibria this family appeared in work by O’Neil and Cox-Steib [27] (also see [18]). There, sequences start from $P_{0}=1$, $P_{1}=z^{1/2}$. In difference from the Adler-Moser case, these sequences terminate as a non-factorizable $\varkappa$ emerges at some step of recurrence. Examples of first several sequences are:
@@
Similar situation occurs in the case of the third-order operators. As we saw in section 5, the Darboux transformations (31, 36), together with (37), result in the first-order recurrence relations (8). For two species of charges initial conditions are $p_{0}=q_{0}=1$. However, there exist other initial conditions, where terminating sequences of equilibrium configurations, that are not solutions of (7), are generated by the Darboux transformations. Such sequences of solutions were found in [27] (also see [10], [18]). For example, for $q_{0}=1$, $p_{0}=z^{2}$ the first several sequences are
@@ 7 Conclusions and Open Problems
When $\Lambda=1$, this generalization is associated with the soliton solutions of the KdV hierarchy and “trigonometric” Baker-Akhieser function $\Phi=p(z,k)/q(z)e^{kz}$, where now $q$ and $p$ are trigonometric polynomials in $z$ ([14, 17, 21], for review see e.g. [22]). Trigonometric Baker-Akhiezer function is obtained from the eigenfunction $e^{kz}$ (with eigenvalue $\lambda=k^{2}$) of the free Schrodinger operator $\partial^{2}$ by iterations of the Darboux transformations: First with $\eta=k_{1}^{2}$, then transformation with $\eta=k_{2}^{2}$, etc. Here $k_{i}$ are integers, such that $0<k_{1}<k_{2}<\dots$.
When $\Lambda=2$, translating vortex streets configurations cannot be generated by Darboux transformations: Even in the non-translating case $k=0$, only short terminating sequences are found [22]. This is because the operator $L$ is now of the third order. Moreover the “free” operator $L_{0}=-\partial^{3}$ does not have trigonometric eigenfunctions and one cannot expect to find non-terminating sequences here. It would be interesting to explore the $\Lambda=2$ vortex street configurations using alternative methods.